Kp for N2O4<------>2NO2 at equilibrium pressure 3Po is 0.5Po on halving the volume of container equilibrium is distributed the new equilibrium pressure will be?
1
Expert's answer
2019-05-08T04:34:52-0400

N2O4 = 2 NO2

Kp = 0.5^2/3 = 0.083

P = 1/V

P2 = 2*P1